OpenStudy (anonymous):

I need to know how to solve 5/13x=20

OpenStudy (zzr0ck3r):

\[\frac{5}{13x}=20\]?

OpenStudy (anonymous):

Yes that's the problem. If you can help me out sure will appreciate it

OpenStudy (zzr0ck3r):

multiply both sides by x \[x*\frac{5}{13x}=20*x\] the x's on the left cancel and we are left with \[\frac{5}{13}=20x\] now divide each side by 20 \[\frac{5}{13*20}=x\]

OpenStudy (anonymous):

Thank you very much

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

How do I solve m=8/9,(2,-1)